# EDADepth: Enhanced Data Augmentation for Monocular Depth Estimation

### v1

@@ Abstract
Due to their text-to-image synthesis feature, diffusion models have recently seen a rise in visual perception tasks, such as depth estimation. The lack of good-quality datasets makes the extraction of a fine-grain semantic context challenging for the diffusion models. The semantic context with fewer details further worsens the process of creating effective text embeddings that will be used as input for diffusion models. In this paper, we propose a novel EDADepth, an enhanced data augmentation method to estimate monocular depth without using additional training data. We use Swin2SR, a super-resolution model, to enhance the quality of input images. We employ the BEiT pre-trained semantic segmentation model for better extraction of text embeddings. We introduce BLIP-2 tokenizer to generate tokens from these text embeddings. The novelty of our approach is the introduction of Swin2SR, the BEiT model, and the BLIP-2 tokenizer in the diffusion-based pipeline for the estimation of monocular depth. Our model achieves state-of-the-art results (SOTA) on the $\delta_{3}$ metric on both NYUv2 and KITTI datasets. It also achieves results comparable to those of the SOTA models in the RMSE and REL metrics. Finally, we also show improvements in the visualization of the estimated depth in comparison to the SOTA diffusion-based monocular depth estimation models. Anonymous code repository: https://github.com/edadepthmde/EDADepth_ICMLA.
@@
Monocular Depth Estimation, Semantic Context, Text Embbedings, Tokenizer.
@@ I Introduction
Depth estimation is an essential task in computer vision that measures the distance of each pixel relative to a camera. Depth is necessary for operations such as 3D reconstruction [1] (refer 1) and scene understanding [2]. One of the types of depth estimation is Monocular Depth Estimation (MDE) [3]. MDE is a task that estimates the depth of an object using a single-view image. Since single-view images do not have epipolar geometry [4], it is challenging to determine the depth of each pixel. Traditional methods for depth estimation used monocular cues [5] and shading [6]. However, such methods faced challenges, such as the varying image lighting and the need for precise camera calibration. Such limitations suggested a need for a technique to estimate depth value based on per-pixel regression[7], a task commonly used in deep learning  [3]. Hence, deep learning methods have emerged as a reliable solution for depth estimation.

### v2

@@ Abstract
Due to their text-to-image synthesis feature, diffusion models have recently seen a rise in visual perception tasks, such as depth estimation. The lack of good-quality datasets makes the extraction of a fine-grain semantic context challenging for the diffusion models. The semantic context with fewer details further worsens the process of creating effective text embeddings that will be used as input for diffusion models. In this paper, we propose a novel EDADepth, an enhanced data augmentation method to estimate monocular depth without using additional training data. We use Swin2SR, a super-resolution model, to enhance the quality of input images. We employ the BEiT pre-trained semantic segmentation model to better extract text embeddings. We use the BLIP-2 tokenizer to generate tokens from these text embeddings. The novelty of our approach is the introduction of Swin2SR, the BEiT model, and the BLIP-2 tokenizer in the diffusion-based pipeline for the monocular depth estimation. Our model achieves state-of-the-art results (SOTA) on the $\delta_{3}$ metric on the NYUv2 and KITTI datasets. It also achieves results comparable to those of the SOTA models in the RMSE and REL metrics. Finally, we also show improvements in the visualization of the estimated depth compared to the SOTA diffusion-based monocular depth estimation models—code repository: https://github.com/edadepthmde/EDADepth_ICMLA.
@@
Monocular Depth Estimation, Semantic Context, Text Embeddings, Tokenizer.
@@ I Introduction
Depth estimation is an essential task in computer vision that measures the distance of each pixel relative to a camera. Depth is necessary for operations such as 3D reconstruction [1] (refer 1), scene understanding [2] and recognition [3]. One of the types of depth estimation is Monocular Depth Estimation (MDE) [4]. MDE is a task that estimates the depth of an object using a single-view image. Since single-view images do not have epipolar geometry [5], it is challenging to determine the depth of each pixel. Traditional methods for depth estimation used monocular cues [6] and shading [7]. However, such methods faced challenges, such as the varying image lighting and the need for precise camera calibration. Such limitations suggested a need for a technique to estimate depth value based on per-pixel regression[8], a task commonly used in deep learning  [4]. Hence, deep learning methods have emerged as a reliable solution for depth estimation.